OpenStudy (anonymous):

Two numbers have prime factorizations of 2 2 · 3 · 5 and 2 · 3 2 · 7. Which expression can be used to find their least common multiple?

OpenStudy (anonymous):

Is this correct? 22 * 3 * 5 2 * 32 * 7

OpenStudy (anonymous):

I believe you mean that the first one is \[2^2\cdot3\cdot5\] and the second is \[2\cdot 3^2 \cdot 7\] In which case, the least common multiple is found by taking the max power of each term. Thus, the lcm is \[2^2 \cdot 3^2 \cdot 5 \cdot 7 = 1260\]

OpenStudy (anonymous):

is kenz in the house?